OpenStudy (loser66):

Evaporation of sweat is the human body's cooling mechanism. At body temperature, it takes 2.4 MJ/kg to evaporate water. Marathon runners typically lose about 3 L of sweat each hour.How much energy gets lost to sweating during a 1.7-hour marathon?

OpenStudy (loser66):

This is my work: total mass of sweat: 3*1.7 = 5.1L = 5.1kg Heat-loss : Q = mL = 5.1 kg* 2.4*10^6 J/kg = 1.224*10^8 J and then, energy P = Q/t = 7.2 *10^6 J/s Am I right? @ybarrap

OpenStudy (loser66):

oh, no, P = Q/t = 1.224*10^8 J/ 6120s = 20000J/s= 20kJ/s

OpenStudy (ybarrap):

The only problem I see is that P is power not energy. Q is what you are looking for.

OpenStudy (agent0smith):

"'How much energy gets lost " You already found that. You went too far. Energy is a quantity in joules, NOT joules per second... that'd be a rate of energy usage/production.

OpenStudy (loser66):

:) so, it's just Q? how about P in P = Q/t ? what is it?

OpenStudy (agent0smith):

Power = energy/time

OpenStudy (ybarrap):

That is power, rate of change of energy per unit time

OpenStudy (loser66):

Thanks you all. :)

OpenStudy (ybarrap):

That is very easy to confuse. You walk a mile, you run a mile. You use the same amount of energy. But power is different in each case.

OpenStudy (loser66):

because the time used in walking is longer than it is in running, right?

OpenStudy (agent0smith):

Yes so power will be lower when walking.

OpenStudy (loser66):

Thank you. :)